OpenStudy (anonymous):

what is the point of intersection for y=3x+1 and 4x-2y=10

OpenStudy (precal):

solve for x or y

OpenStudy (precal):

y=3x+1 4x-2y=10 y=(10-4x)/-2

OpenStudy (precal):

y=-5+2x

OpenStudy (precal):

set them both equal to each other and solve for x 3x+1=-5+2x

OpenStudy (precal):

once you know your x value then sub x into one of the equations and solve for y

OpenStudy (anonymous):

how did you get y=-5x+2

OpenStudy (precal):

i took your second equation and solved for y

OpenStudy (precal):

-4x to both sides and divide everything by -2

OpenStudy (anonymous):

i got y=-3x because when you subtract the 4x you would get -2y=6x which would then equal -3x

OpenStudy (anonymous):

i see where i messed up. Thank you

OpenStudy (precal):

you are welcome